A. Assuming a two-terminal circuit element with impedance Z = 2.3 + 5i is driven by a sinusoidal current I = π’Š(𝟎.πŸ’πŸβˆ’πŸ‘.πŸ”π’Š) πŸ‘.πŸ“+𝟐.πŸ“π’Š . Find the voltage of the circuit (according to Ohm’s law V = IZ).Β 


1
Expert's answer
2021-09-08T05:26:35-0400

Formula, V=IZ

"\implies V =[i(0.42-3.6i)3.5+2.5i].(2.3+5i)\\
=(12.6+3.97i).(2.3+5i)\\
=12.6×2.3+3.97i×2.3+12.6×5i+3.97i×5i\\
=9.13+72.131i\\
\text{Thus ,the required voltage is 9.13+72.131i.}"